OpenStudy (anonymous):

What are the directrices of the ellipse given by the equation x^2/36+y^2/100=1

OpenStudy (anonymous):

\[\frac{ x^2 }{ 36 }+\frac{ y^2 }{ 100 }=1\]

OpenStudy (anonymous):

Choices: x= +-5/8 y= +-5/8 x= +-25/2 y= +-25/2

OpenStudy (jdoe0001):

directrix of an ellipse will be at \(\large \cfrac{a^2}{c}\) distance from the center of the ellipse

OpenStudy (jdoe0001):

so, which one above would be your "a" ?

OpenStudy (jdoe0001):

bearing in mind that the "c" distance from the center to a focus is \(\large c^2 = a^2-b^2\)

OpenStudy (anonymous):

36?

OpenStudy (jdoe0001):

for an ellipse the "a" element is the "bigger one" of the 2 denominators, and at the "axis" is where the ellipse has its major axis

OpenStudy (anonymous):

So a would be 100

OpenStudy (jdoe0001):

right, so what would you get for "c" off \(\large c=\sqrt{a^2-b^2}\)

OpenStudy (anonymous):

1? Not sure.

OpenStudy (jdoe0001):

well, your elipse equation is $$ \cfrac{ x^2 }{ \color{red}{36} }+\cfrac{ y^2 }{ \color{red}{100} }=1 \implies \cfrac{ (x-h)^2 }{ \color{red}{b^2} }+\cfrac{ (y-k)^2 }{ \color{red}{a^2} }=1 $$

OpenStudy (jdoe0001):

keep in mind that 36 and 100 are in squared form if you were to write them without the squared version, they'd be $$ \cfrac{ x^2 }{ \color{red}{6^2} }+\cfrac{ y^2 }{ \color{red}{10^2} }=1 \implies \cfrac{ (x-h)^2 }{ \color{red}{b^2} }+\cfrac{ (y-k)^2 }{ \color{red}{a^2} }=1 $$